OpenStudy (anonymous):

What is the sum of a 6–term geometric series if the first term is 21 and the last term is 1,240,029?

OpenStudy (anonymous):

common ratio: ar^5 / a = r^5 so r^5 = 1240029 / 21 = 59049 r = 9

OpenStudy (anonymous):

sum of 6 terms = 21 * (9^6 - 1) ------- 6 - 1 = 2232048